# Kaon Modification in Pion Medium

### v1

@@ Abstract
Kaon properties in pion medium are studied. The dense pion medium is produced in high-energy heavy-ion collisions. The consideration is based on the chiral nucleon-kaon-pion Lagrangian. It is found that the pion medium does not produce a substantive contribution but enhances the effect of the baryon matter. Numerical estimate shows that this enhancement factor induced by the pion medium does not exceed 11% at the freeze-out stage of heavy-ion collisions, i.e. it is small. However, the impact of this enhancement can be higher in actual nuclear collisions because the effect of the in-medium (anti)kaon modification is accumulated during the (anti)kaon evolution before the freeze-out when the pion density is higher.
@@
In all aforementioned studies the in-medium modifications of kaons were induced by the baryon medium. At higher collision energies, mesons (mostly pions) are abundantly produced. In this paper we address the question if the pionic medium can modify the kaon properties additionally to the baryon-induced modifications?
The present consideration follows the lines of Ref. Kolomeitsev:1995we , where the influence of the Bose-Einstein pion condensate on kaon properties was studied. In the present paper, the effect of the thermal pion medium on kaon properties is considered.
@@
$f_{\pi}=$ 93 MeV is the pion decay constant, $f_{K}$ is the kaon decay constant, and $\Sigma_{KN}$ is the kaon-nucleon sigma term Hartnack:2011cn . The term proportional to $\Sigma_{KN}$ results from the attractive scalar interaction due to explicit chiral symmetry breaking. The $\pi N$ interaction is omitted in this Lagrangian because it is irrelevant to our purposes. The chiral $\pi K$ Lagrangian contains additional $1/2$ factor as compared to Refs. Kolomeitsev:1995we ; Belkov:1988gg because scalar $\boldsymbol{\pi}$ fields are used here instead of charged $\pi^{+}$ and $\pi^{-}$ in Refs. Kolomeitsev:1995we ; Belkov:1988gg .
@@
where the upper (lower) sign in front of the $\omega\rho_{N}$ term corresponds to the (anti)kaon and $(\omega,{\bf k})$ is the four-momentum of the (anti)kaon, The quantity
$m_{\pi}\langle\boldsymbol{\pi}^{2}\rangle=\rho_{S}^{\pi}=3m_{\pi}\int\frac{d^{3}k}{(2\pi)^{3}\omega_{\pi}(k)}\frac{1}{\exp(\omega_{\pi}/T)-1}$ (16)
is the scalar pion density with $T$ being the temperature of the medium.
@@
The result of numerical calculation of the enhancement factor $F_{\pi}$ induced by the pion medium at the freeze-out stage as a function the center-of-mass energy of the nuclear collision is presented in Fig. 1. As seen, it does not exceed 11%, i.e. it is small. This retroactively confirms the approximations made at derivation of Eq. (17). However, the impact of this enhancement can be higher in actual nuclear collisions because the effect of the in-medium (anti)kaon modification is accumulated during the (anti)kaon evolution before the freeze-out when the pion density is higher.
Enlightening discussions with D.N. Voskresensky are gratefully acknowledged. This work was carried out using computing resources of the federal collective usage center “Complex for simulation and data processing for mega-science facilities” at NRC ”Kurchatov Institute” ckp.nrcki.ru .

### v2

@@ Abstract
Kaon properties in thermal pion-nucleon medium are studied. The dense pion-nucleon medium is produced in high-energy heavy-ion collisions. The consideration is based on the chiral nucleon-kaon-pion Lagrangian. It is found that the pion medium does not produce a substantive contribution but enhances the effect of the baryon matter. Numerical estimate shows that this enhancement factor induced by the pion medium does not exceed 5% at the freeze-out stage of heavy-ion collisions, i.e. it is small. However, the impact of this enhancement can be higher in actual nuclear collisions because the effect of the in-medium (anti)kaon modification is accumulated during the (anti)kaon evolution before the freeze-out when the pion density is higher.
@@
In all aforementioned studies the in-medium modifications of kaons were induced by the baryon medium. At higher collision energies, mesons (mostly pions) are abundantly produced. Therefore, we address the question if the pionic medium can modify the kaon properties additionally to the baryon-induced modifications? This question has been already studied in Ref. Faessler:2002qb , where the kaons were considered in purely pionic medium. It was found that the main effect of that purely pionic medium is the kaon broadening due to rescatterings inside of the pion matter. In particular, this broadening results in an acceleration of the $\phi\to K^{+}K^{-}$ decays. In the present paper the kaon modification is studied in combined pion-nucleon medium.
The present consideration follows the lines of Ref. Kolomeitsev:1995we , where the influence of the Bose-Einstein pion condensate (with nonzero pion chemical potential) on kaon properties was studied. In the present paper, the effect of the thermal pion-nucleon medium (with zero pion chemical potential) is considered.
@@
$f_{\pi}=$ 93 MeV is the pion decay constant, $f_{K}$ is the kaon decay constant that approximately equals $f_{\pi}$, and $\Sigma_{KN}$ is the kaon-nucleon sigma term Reya:1974gk . The term proportional to $\Sigma_{KN}$ results from the attractive scalar interaction due to explicit chiral symmetry breaking. The numerical values of the $\Sigma_{KN}$ term, derived from experimental data, vary widely Reya:1974gk . In practical calculations, the value of 350 MeV is usually used Hartnack:2011cn ; Ko:1996yy . The $\pi N$ interaction is omitted in this Lagrangian because it is irrelevant to our purposes. The chiral $\pi K$ Lagrangian contains additional $1/2$ factor as compared to Refs. Kolomeitsev:1995we ; Belkov:1988gg because scalar $\boldsymbol{\pi}$ fields are used here instead of charged $\pi^{+}$ and $\pi^{-}$ in Refs. Kolomeitsev:1995we ; Belkov:1988gg .
@@
where the upper (lower) sign in front of the $\omega\rho_{N}$ term corresponds to the (anti)kaon and $(\omega,{\bf k})$ is the four-momentum of the (anti)kaon, The quantity $\langle\boldsymbol{\pi}^{2}\rangle$ reads
$\langle\boldsymbol{\pi}^{2}\rangle=\frac{\rho_{S}^{\pi}}{2m_{\pi}}=g_{\pi}\int\frac{d^{3}k}{(2\pi)^{3}2\omega_{\pi}(k)}\frac{1}{\exp(\omega_{\pi}/T)-1}$ (16)
where $\rho_{S}^{\pi}$ is the scalar pion density, $g_{\pi}=3$ is the pion degeneracy factor, and $T$ is the temperature of the medium.
@@
The result of numerical calculation of the enhancement factor $F_{\pi}$ induced by the pion medium at the freeze-out stage as a function the center-of-mass energy of the nuclear collision is presented in Fig. 1. As seen, it does not exceed 5%, i.e. it is small. This retroactively confirms the approximations made at derivation of Eq. (17). However, the impact of this enhancement can be higher in actual nuclear collisions because the effect of the in-medium (anti)kaon modification is accumulated during the (anti)kaon evolution before the freeze-out when the pion density is higher.
Enlightening discussions with D.N. Voskresensky and E.E. Kolomeitsev are gratefully acknowledged. This work was carried out using computing resources of the federal collective usage center “Complex for simulation and data processing for mega-science facilities” at NRC ”Kurchatov Institute” ckp.nrcki.ru .